[Brahms-l] Shift report 20040229 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Sun Feb 29 2004 - 08:12:26 EST
------------------------------------------------------------
Date: Sun Feb 29 08:12:26 2004, Shift: 00:00-08:00
Supervisor: rw, with: 
Log book#15 pages 70 - 70

Shift Summary:
------------------------------------------------------------
1. Beam is on from the previous shift to 0:30, from 1:00 to 5:00 and from
   5:40 to the end of the shift. A very nice shift.

2.  run      tr2      tr3
   10667     22k      55k     FS 8deg@1723B   MRS 40deg.@2255B
   10668    TPC Pedestal
   10669    Non-TPC Pedestal
   10670     81k      99k
   10671     88k     136k
   10672     62k     117k
   10673     50K     125k
   10674     12k      50k     MRS 10M events completed.     
   10675    TPC Pedestal 
   10676    Non-TPC pedestal
   10677     92k     202k     MRS 40deg.@1050B   FS are as before
   10679     50k     110k
   10680     running

3, no detector problem. 
 
4. DAQ run error occured and can't start a new run.    
   Isoserver and run server restarted. The error resumed. 
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Feb 29 08:13:16 2004

This archive was generated by hypermail 2.1.8 : Sun Feb 29 2004 - 08:13:31 EST